A 0.050 0-kg ingot of metal is heated to 200.0 deg
Answer #1

in the equlibrium condition we use energy conservation

heat released from the heated metal=heat is absorbed by water

m_{metal}C_{metal}\Delta \Theta _{metal}=m_{water}C_{water}\Delta \Theta _{water}

0.05*C_{metal}(200-22.4)=0.4*4186*(22.4-20)

C_{metal}=452.54J/Kg^{\circ}C

heat transfer to the water H=mC\Delta \Theta

  H=0.4*4186*(22.4-20)

H=4018.56 J